# CoVR: Learning Composed Video Retrieval from Web Video Captions

### v3

@@ Abstract
Composed Image Retrieval (CoIR) has recently gained popularity as a task that considers both text and image queries together, to search for relevant images in a database. Most CoIR approaches require manually annotated datasets, comprising image-text-image triplets, where the text describes a modification from the query image to the target image. However, manual curation of CoIR triplets is expensive and prevents scalability. In this work, we instead propose a scalable automatic dataset creation methodology that generates triplets given video-caption pairs, while also expanding the scope of the task to include composed video retrieval (CoVR). To this end, we mine paired videos with a similar caption from a large database, and leverage a large language model to generate the corresponding modification text. Applying this methodology to the extensive WebVid2M collection, we automatically construct our WebVid-CoVR dataset, resulting in 1.6 million triplets. Moreover, we introduce a new benchmark for CoVR with a manually annotated evaluation set, along with baseline results. Our experiments further demonstrate that training a CoVR model on our dataset effectively transfers to CoIR, leading to improved state-of-the-art performance in the zero-shot and finetuning setups on both the CIRR and FashionIQ benchmarks. Our code, datasets, and models are publicly available at https://imagine.enpc.fr/~ventural/covr.
@@ 1 Introduction
CoVR has many use cases, including but not limited to searching online videos for finding reviews of a specific product, how-to videos of a tool for specific usages, live events in specific locations, sports matches of specific players. Similar to composed image retrieval (CoIR), CoVR is also particularly useful when conveying a concept with a visual is easier and/or more accurate than only using words (e.g., unknown location/object, specific camera view, specific color).
@@
Given the increased momentum in vision and language research in the recent years (Li et al. 2022b; Radford et al. 2021), CoIR has emerged as a new task (Vo et al. 2019), and since then witnessed improvements of both models and benchmarks (Baldrati et al. 2023, 2022; Gu et al. 2023; Levy et al. 2024; Liu et al. 2021; Wu et al. 2021). However, to the best of our knowledge, CoVR was not studied before. A key challenge in building CoVR models is the difficulty of gathering suitable training data of video-text-video triplets. We overcome this limitation by developing an automatic approach to generate triplets from existing video-caption collections. Specifically, we mine video pairs whose corresponding captions slightly differ in text space. We automatically describe this difference with a language model, which we train for a modification-text generation task. In particular, we use manually annotated triplets, each containing: (a) source caption, (b) target caption, (c) the modification text. We then finetune a large language model (LLM) (Touvron et al. 2023) by inputting (a-b), and outputting (c). We assume the resulting modification to describe the difference between the corresponding videos, thus obtaining video-text-video triplets (see Figure 2 for an overview). When training our CoVR/CoIR models, we can flexibly select one or more frames from the videos, enabling multiple settings (i.e., retrieving images or videos).
@@
Figure 2: Method overview: We automatically mine similar caption pairs from a large video-caption database from the Web, and use our modification text generation language model (MTG-LLM) to describe the difference between the two captions. MTG-LLM is trained on a dataset of 715 triplet text annotations (Brooks, Holynski, and Efros 2023). The resulting triplet with the two corresponding videos (query $q$ and target video $v$) and the modification text ($t$) is therefore obtained fully automatically, allowing a scalable CoVR training data generation.
@@
We apply our triplet generation approach to the WebVid2M dataset (Bain et al. 2021) which contains 2.5M Web-scraped video-caption pairs. This results in the WebVid-CoVR training dataset with 1.6M CoVR triplets. By virtue of its automatic generation procedure, WebVid-CoVR is inherently noisy. To efficiently train on such large-scale and noisy data, we use a contrastive loss (van den Oord, Li, and Vinyals 2018), adopting the HN-NCE variant from (Radenovic et al. 2023) to upsample the significance of hard negatives. We design a CoVR model based on the cross-modal BLIP (Li et al. 2022b) and use query scoring (Bain et al. 2022) to exploit information from multiple video frames. Training this model on WebVid-CoVR shows strong transferability to the CoIR task, in both zero-shot and finetuning settings, achieving state-of-the-art results on the standard CIRR (Liu et al. 2021) and FashionIQ (Wu et al. 2021) benchmarks. Finally, to foster research in CoVR, we repeat our generation procedure on a distinct subset of the WebVid10M dataset (Bain et al. 2021) and manually select correctly generated samples to constitute WebVid-CoVR-Test, a test set of 2,435 CoVR triplets. We find that our model achieves promising results on WebVid-CoVR-Test compared to standard baselines.
@@
To summarize, our contributions are: (i) We propose a scalable approach to automatically generate composed visual retrieval training data. We apply this pipeline to the WebVid2M dataset and generate the WebVid-CoVR training dataset with 1.6M CoVR triplets. (ii) We show that training a CoVR model on WebVid-CoVR transfers well to the CoIR task, and achieves state-of-the-art results on the CIRR and FashionIQ benchmarks in the zero-shot and finetuning setups. (iii) We evaluate our model on WebVid-CoVR-Test, a new CoVR benchmark that we manually annotate. Our code, datasets, and models are available at https://imagine.enpc.fr/~ventural/covr.
@@ 2 Related Work
Dataset Type #Triplets #Visuals #Unique words Avg. text length Domain
@@
Composed image retrieval (CoIR). CoIR (Vo et al. 2019) has been an active area of research in recent years (Guo et al. 2018; Vo et al. 2019; Wu et al. 2021; Baldrati et al. 2022; Liu et al. 2021; Delmas et al. 2022; Kim et al. 2021; Baldrati et al. 2023; Saito et al. 2023; Gu et al. 2023). Most methods designed for this problem use manually annotated image-text-image triplets for training (Wu et al. 2021; Liu et al. 2021; Baldrati et al. 2022; Delmas et al. 2022). Very recent works, such as Pic2Word (Saito et al. 2023) and SEARLE (Baldrati et al. 2023), explore zero-shot CoIR setups where no manually annotated CoIR triplet is used. These approaches build on CLIP (Radford et al. 2021) and train a mapping network using image-only data for text inversion so that they can be flexibly composed with text descriptions. Our approach is similar in that it avoids collecting manual triplets; however, we instead perform supervised training on automatically generated image-text-video triplets given only video-text pairs. We also differ from above works by focusing on the composed video retrieval (CoVR) task, as opposed to CoIR.
@@
Datasets for composed image retrieval. CIRR (Liu et al. 2021) and Fashion-IQ (Wu et al. 2021) are the two most widely used CoIR benchmarks. Both are manually annotated, hence small scale (about 30k triplets, see Table 1) due to the high cost implied in collecting CoIR triplets. To scale up, two concurrent works proposed larger, automatically generated CoIR datasets: LaSCo (Levy et al. 2024) and SynthTriplets18M (Gu et al. 2023). However, these two datasets are currently not publicly available. The LaSCo dataset (Levy et al. 2024) is generated using the visual question answering annotations and the pairing between images and counterfactual images in the VQAv2 dataset (Antol et al. 2015). In detail, this dataset provides for each (image, question, answer) triplet a counterfactual triplet with the same question and different image and answer. In contrast, we do not rely on such expensive annotation schemes. SynthTriplets18M (Gu et al. 2023) uses the text-conditioned image editing framework InstructPix2Pix (Brooks, Holynski, and Efros 2023) to automatically generate CoIR data. Their edit text generation process is similar to ours, but our generation process differs in that we automatically mine similar videos from a dataset of video-text pairs to construct CoVR triplets instead of generating visual data. In experiments, we show the superiority of our triplet construction procedure as we achieve much higher CoIR results (e.g., 39% vs 19% zero-shot R@1 on CIRR while generating fewer data). Lastly, our WebVid-CoVR dataset is not limited to still images and considers videos, while standing out as the largest composed retrieval dataset in the natural domain, as depicted in Table 1.
@@
Vision-language pretraining. Many strong multi-modal models have been pretrained on large datasets of image-caption pairs (Alayrac et al. 2022; Chen et al. 2020; Li et al. 2021, 2023, 2020b; Radford et al. 2021; Jia et al. 2021; Su et al. 2019; Schuhmann et al. 2022) or video-caption pairs (Akbari et al. 2021; Li et al. 2022a, 2020a; Miech et al. 2019, 2020; Xu et al. 2021; Sun et al. 2022; Xue et al. 2022a; Zhao et al. 2023). In contrast, we generate CoVR training data from video-caption pairs instead of directly training on them. Our data generation approach is also related to other generation approaches used for other tasks, e.g., action recognition (Nagrani et al. 2020), visual question answering (Yang et al. 2021), and visual dialog (Liu et al. 2023). However, unlike all these tasks, the CoVR task requires retrieving visual data.
@@
Video retrieval. Text-to-video retrieval has received great attention over the last few years (Fang et al. 2021; Gao et al. 2021; Ge et al. 2022; Liu et al. 2022; Luo et al. 2021; Ma et al. 2022; Rasheed et al. 2023; Xue et al. 2022b; Yang, Bisk, and Gao 2021; Yao et al. 2022; Xu et al. 2021). We also make use of multiple video frames with query scoring similar to (Bain et al. 2022). However, different from these methods, we focus on composed video retrieval, where the query consists of both text and visual data.
@@ 3 Automatic Triplet Generation and Training
The goal of our composed video retrieval (CoVR) task is, given an input image $q$ and a modification text $t$, to retrieve a modified video $v$ in a large database of videos [^1] Note that $q$ could also be a video query, but in our main experiments we focus on an image query, and provide more results in the supplementary material (Section C.2) with video queries.. Our goal is to avoid the manual annotation of $(q,t,v$) triplets for training. Hence we automatically generate such triplets from Web-scraped video-caption pairs, as explained in Section 3.1 and illustrated in Figure 2. The resulting WebVid-CoVR dataset, together with its manually curated evaluation set, is presented in Section 3.2. Finally, we present how we train a CoVR model using WebVid-CoVR in Section 3.3.
@@ 3.1 Generating Composed Video Retrieval Triplets
Given a large (Web-scraped) dataset of video-caption pairs, we wish to automatically generate video-text-video CoVR triplets $(q,t,v$) where the text $t$ describes a modification to the visual query $q$. However, the dataset of video-caption pairs neither contains annotations of paired videos, nor modification text that describes their difference. Hence we propose a methodology to automatically mine paired videos and describe their difference, as described below. Note that for illustration, we take as an example the WebVid2M dataset (Bain et al. 2021) with 2.5M video-caption pairs, but this methodology could potentially be applied to other large datasets of video-text (or image-text) pairs.
@@
Filtering caption pairs. We wish to automatically generate the modification text between paired videos using their (paired) captions. However, caption pairs with the same meaning are likely to result in meaningless differences. On the contrary, caption pairs that differ too much are likely to result in large visual differences that cannot be easily described. To address these issues, we filter out caption pairs that are too similar and too dissimilar. Specifically, we exclude caption pairs with CLIP text embedding similarity $\geq$ 0.96 (e.g., “Fit and happy young couple playing in the park” and “Fit and happy young couple play in the park”) and caption pairs with CLIP text embedding similarity $\leq$ 0.6 (e.g., Zebra on a white background” and Coins on a white background”). We also exclude pairs where the captions differ by a digit (which mostly consist of a date in practice), a word not part of the English dictionary, or by a rare word. Rare words are detected based on the zipfzipf frequency (Speer 2022). Finally, we remove templated captions such as “abstract of”, “concept of”, and “flag of” which are over-represented in WebVid2M. At the end of this filtering stage, we have 370k distinct caption pairs with 92k distinct captions, resulting in 1.2M paired videos that we will use to generate the modification text. Note that we can use these paired videos in both directions to generate triplets, as the source and target videos can be swapped.
@@
Generating a modification text from paired captions. In order to generate a modification text between paired videos, we develop and apply a “modification text generation large language model” (MTG-LLM) to their corresponding paired captions. We describe the MTG-LLM inference process below and then explain its training details. The MTG-LLM takes as input two paired captions and generates a modification text that describes the difference between the two captions (see Figure 2). In detail, the generation is auto-regressive, i.e., we recursively sample from the token likelihood distribution conditioned on the previously generated tokens until an end-of-sentence token is reached. Examples of the input-output, and details about the prompt format, which involves concatenating the two captions with a delimiter, can be found in Section B.4 of the Appendix. We use top-k sampling (Fan, Lewis, and Dauphin 2018) for generating the tokens instead of maximum-likelihood methods such as beam search. Note that we only generate a single modification text per caption pair for computational efficiency, but the MTG-LLM could be used to generate multiple modification texts per caption pair which could serve as a data augmentation in future work.
@@
We now describe the training details of the MTG-LLM. We start from a LLM pretrained with a next token prediction objective on a Web-scale text dataset, namely LLaMA (Touvron et al. 2023). We then finetune this LLM for the MTG task on a manually annotated text dataset. In particular, we repurpose the editing dataset from InstructPix2Pix (Brooks, Holynski, and Efros 2023), which provides a modification text and a target caption for 700 input captions. We augment this dataset with 15 annotations that cover additional cases. More details about the additional examples can be found in Section B.4 of the Appendix.

### v4

# CoVR-2: Automatic Data Construction for Composed Video Retrieval

@@ Abstract
Composed Image Retrieval (CoIR) has recently gained popularity as a task that considers both text and image queries together, to search for relevant images in a database. Most CoIR approaches require manually annotated datasets, comprising image-text-image triplets, where the text describes a modification from the query image to the target image. However, manual curation of CoIR triplets is expensive and prevents scalability. In this work, we instead propose a scalable automatic dataset creation methodology that generates triplets given video-caption pairs, while also expanding the scope of the task to include Composed Video Retrieval (CoVR). To this end, we mine paired videos with a similar caption from a large database, and leverage a large language model to generate the corresponding modification text. Applying this methodology to the extensive WebVid2M collection, we automatically construct our WebVid-CoVR dataset, resulting in 1.6 million triplets. Moreover, we introduce a new benchmark for CoVR with a manually annotated evaluation set, along with baseline results. We further validate that our methodology is equally applicable to image-caption pairs, by generating 3.3 million CoIR training triplets using the Conceptual Captions dataset. Our model builds on BLIP-2 pretraining, adapting it to composed video (or image) retrieval, and incorporates an additional caption retrieval loss to exploit extra supervision beyond the triplet, which is possible since captions are readily available for our training data by design. We provide extensive ablations to analyze the design choices on our new CoVR benchmark. Our experiments also demonstrate that training a CoVR model on our datasets effectively transfers to CoIR, leading to improved state-of-the-art performance in the zero-shot setup on the CIRR, FashionIQ, and CIRCO benchmarks. Our code, datasets, and models are publicly available at https://imagine.enpc.fr/~ventural/covr.
Index Terms:
Composed Video Retrieval, Composed Image Retrieval.
@@ 1 Introduction
CoVR has many use cases, including but not limited to searching online videos for finding reviews of a specific product, how-to videos of a tool for specific usages, live events in specific locations, sports matches of specific players. Similar to composed image retrieval (CoIR), CoVR is also particularly useful when conveying a concept with a visual is easier and/or more accurate than only using words (e.g., unknown location/object, a specific camera view, a specific color).
@@
Given the increased momentum in vision and language research in the recent years [1, 2], CoIR has emerged as a new task [3], and since then witnessed improvements of both models and benchmarks [4, 5, 6, 7, 8, 9]. However, to the best of our knowledge, CoVR was not studied before. A key challenge in building CoVR models is the difficulty of gathering suitable training data of video-text-video triplets. We overcome this limitation by developing an automatic approach to generate triplets from existing video-caption collections. Specifically, we mine video pairs whose corresponding captions slightly differ in text space. We automatically describe this difference with a language model, which we train for a modification-text generation task. In particular, we use manually annotated triplets, each containing: (a) source caption, (b) target caption, (c) the modification text. We then finetune a large language model (LLM) [10] by inputting (a-b), and outputting (c). We assume the resulting modification to describe the difference between the corresponding videos, thus obtaining video-text-video triplets (see Figure 2 for an overview). When training our CoVR/CoIR models, we can flexibly select one or more frames from the videos, enabling multiple settings (i.e., retrieving images or videos).
@@
Figure 2: Method overview: We automatically mine similar caption pairs from a large video-caption database from the Web, and use our modification text generation language model (MTG-LLM) to describe the difference between the two captions. MTG-LLM is trained on a dataset of 715 triplet text annotations [11]. The resulting triplet with the two corresponding videos (query $q$ and target video $v$) and the modification text ($t$) is therefore obtained fully automatically, allowing a scalable CoVR training data generation.
@@
We apply our triplet generation approach to two seed datasets: (1) WebVid2M [12] with 2.5M video-caption pairs, and (2) Conceptual Captions [13] with 3.3M image-caption pairs. We call the resulting training sets as WebVid-CoVR and CC-CoIR, which contain 1.6M CoVR and 3.3M CoIR triplets, respectively. By virtue of its automatic generation procedure, our training data is inherently noisy. To efficiently train on such large-scale and noisy data, we use a contrastive loss [14], adopting the HN-NCE variant from [15] to upsample the significance of hard negatives. Moreover, we integrate an additional contrastive loss, by also retrieving the textual embeddings of the target image, aiming to enhance the attention on the modification text detail. We design a CoVR model based on the cross-modal BLIP-2 [16] and use query scoring [17] to exploit information from multiple video frames. Training this model on WebVid-CoVR shows strong transferability to the CoIR task, in both zero-shot and finetuning settings, achieving state-of-the-art results on the standard CIRR [8], FashionIQ [9], and CIRCO [4] benchmarks in the zero-shot setup. We further improve the performance by jointly training on our WebVid-CoVR and CC-CoIR data together. Finally, to foster research in CoVR, we repeat our generation procedure on a distinct subset of the WebVid10M dataset [12] and manually select correctly generated samples to constitute WebVid-CoVR-Test, a test set of 2,435 CoVR triplets. We find that our model achieves promising results on WebVid-CoVR-Test compared to standard baselines.
@@
To summarize, our contributions are: (i) We propose a scalable approach to automatically generate composed visual retrieval training data. With this methodology, we generate 1.6M WebVid-CoVR and 3.3M CC-CoIR triplets. (ii) We show that training composed retrieval models on our generated datasets transfers well to the CoIR benchmarks, and achieves state-of-the-art results on CIRR, FashionIQ, and CIRCO datasets in the zero-shot setup. (iii) We evaluate our model on WebVid-CoVR-Test, a new CoVR benchmark that we manually annotate. Our code, datasets, and models are publicly available at https://imagine.enpc.fr/~ventural/covr.
@@ 2 Related Work
[…]
Dataset Type #Triplets #Unique visuals #Unique words Avg. text length Domain
@@
Composed image retrieval (CoIR). CoIR [3] has been an active area of research in recent years [18, 3, 9, 5, 8, 19, 20, 4, 21, 6, 22]. Most methods designed for this problem use manually annotated image-text-image triplets for training [9, 8, 5, 19]. Recent works, such as Pic2Word [21], SEARLE [4], TFCIR [23], and LinCIR [24], explore zero-shot CoIR setups where no manually annotated CoIR triplet is used. More recently, Karthik et al. [25] propose a training-free method for CoIR. The approaches [21, 4, 23, 24] build on CLIP [2] and train a mapping network using image-only data for text inversion so that they can be flexibly composed with text descriptions. Our approach is similar in that it avoids collecting manual triplets; however, we instead perform supervised training on automatically generated image-text-video triplets given only video-text pairs. We also differ from above works by focusing on the composed video retrieval (CoVR) task, as opposed to only CoIR.
@@
Datasets for composed image retrieval. CIRR [8] and Fashion-IQ [9] are the two most widely used CoIR benchmarks. Very recently, Baldrati et al. proposed a new test CoIR dataset CIRCO [4], which has gained popularity for having multiple ground truths and many distractors. All three datasets are manually annotated, hence small scale (about 30k triplets, see Table I) due to the high cost implied in collecting CoIR triplets. To scale up, two recent works proposed larger, automatically generated CoIR datasets: LaSCo [7] and SynthTriplets18M [6]. The LaSCo dataset [7] is generated using the visual question answering annotations and the pairing between images and counterfactual images in the VQAv2 dataset [26]. In detail, this dataset provides for each (image, question, answer) triplet a counterfactual triplet with the same question and different image and answer. In contrast, we do not rely on such expensive annotation schemes. SynthTriplets18M [6] uses the text-conditioned image editing framework InstructPix2Pix [11] to automatically generate CoIR data. Their edit text generation process is similar to ours, but our generation process differs in that we automatically mine similar videos from a dataset of video-text pairs to construct CoVR triplets instead of generating visual data. In experiments, we show the superiority of our triplet construction procedure as we achieve much higher CoIR results (e.g., 43.7% vs 26.7% zero-shot R@1 on CIRR while generating fewer data). Similar conclusions hold when pretraining on our automatic CoIR triplets (CC-CoIR). We additionally provide a controlled experiment by training our model on their synthetic images [11, 6] and demonstrate the advantages of using real data. Lastly, our WebVid-CoVR dataset is not limited to still images and considers videos, while standing out as larger than all previous composed retrieval datasets in the natural domain, as depicted in Table I.
@@
Vision-language pretraining. Many strong multi-modal models have been pretrained on large datasets of image-caption pairs [27, 28, 29, 1, 30, 2, 31, 32, 33, 16] or video-caption pairs [34, 35, 36, 37, 38, 39, 40, 41, 42]. In contrast, we generate CoVR training data from video-caption pairs instead of directly training on them. Our data generation approach is also related to other generation approaches used for other tasks, e.g., action recognition [43], visual question answering [44] and visual dialog [45]. However, unlike all these tasks, the CoVR task requires retrieving visual data.
@@
Video retrieval. Text-to-video retrieval has received great attention over the last few years [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 39]. We also make use of multiple video frames with query scoring similar to [17]. However, different from these methods, we focus on composed video retrieval, where the query consists of both text and visual data.
@@ 3 Automatic Triplet Generation and Training
The goal of our composed video retrieval (CoVR) task is, given an input image $q$ and a modification text $t$, to retrieve a modified video $v$ in a large database of videos [^1] Note that $q$ could also be a video query, but in our main experiments we focus on an image query, and provide more results in (Section D.1 of the Appendix) with video queries.. Our goal is to avoid the manual annotation of $(q,t,v$) triplets for training. Hence we automatically generate such triplets from Web-scraped video-caption pairs, as explained in Section 3.1 and illustrated in Figure 2. The resulting WebVid-CoVR dataset, together with its manually curated evaluation set, is presented in Section 3.2. Finally, we present how we train a CoVR model using WebVid-CoVR in Section 3.3.
@@ 3.1 Generating composed video retrieval triplets
Given a large (Web-scraped) dataset of video-caption pairs, we wish to automatically generate video-text-video CoVR triplets $(q,t,v$) where the text $t$ describes a modification to the visual query $q$. However, the dataset of video-caption pairs neither contains annotations of paired videos, nor modification text that describes their difference. Hence we propose a methodology to automatically mine paired videos and describe their difference, as described below. Note that for illustration, we take as an example the WebVid2M dataset [12] with 2.5M video-caption pairs, but this methodology could be applied to other large datasets of video-text (or image-text) pairs. To strengthen our conclusions, we employ the same methodology with the Conceptual Captions image-text dataset [13], which is briefly described along with experiments in Section 4.3. While the rest of this section focuses on CoVR, the data generation pipeline is similar for CoIR.
@@
Filtering caption pairs. We wish to automatically generate the modification text between paired videos using their (paired) captions. However, caption pairs with the same meaning are likely to result in meaningless differences. On the contrary, caption pairs that differ too much are likely to result in large visual differences that cannot be easily described. To address these issues, we filter out caption pairs that are too similar and too dissimilar. Specifically, we exclude caption pairs with CLIP text embedding similarity $\geq$ 0.96 (e.g., “Fit and happy young couple playing in the park” and “Fit and happy young couple play in the park”) and caption pairs with CLIP text embedding similarity $\leq$ 0.6 (e.g., Zebra on a white background” and Coins on a white background”). We also exclude pairs where the captions differ by a digit (which mostly consist of a date in practice), a word not part of the English dictionary, or by a rare word. Rare words are detected based on the zipf frequency [56]. Finally, we remove templated captions such as “abstract of”, “concept of”, and “flag of” which are over-represented in WebVid2M. At the end of this filtering stage, we have 370k distinct caption pairs with 92k distinct captions, resulting in 1.2M paired videos that we will use to generate the modification text. Note that we can use these paired videos in both directions to generate triplets, as the source and target videos can be swapped.
@@
Generating a modification text from paired captions. In order to generate a modification text between paired videos, we develop and apply a “modification text generation large language model” (MTG-LLM) to their corresponding paired captions. We describe the MTG-LLM inference process below and then explain its training details. The MTG-LLM takes as input two paired captions and generates a modification text that describes the difference between the two captions (see Figure 2). In detail, the generation is auto-regressive, i.e., we recursively sample from the token likelihood distribution conditioned on the previously generated tokens until an end-of-sentence token is reached. Examples of the input-output, and details about the prompt format, which involves concatenating the two captions with a delimiter, can be found in Section C.4 of the Appendix. We use top-k sampling [57] for generating the tokens instead of maximum-likelihood methods such as beam search. Note that we only generate a single modification text per caption pair for computational efficiency, but the MTG-LLM could be used to generate multiple modification texts per caption pair which could serve as a data augmentation in future work.
@@
We now describe the training details of the MTG-LLM. We start from a LLM pretrained with a next token prediction objective on a Web-scale text dataset, namely LLaMA [10]. We then finetune this LLM for the MTG task on a manually annotated text dataset. In particular, we repurpose the editing dataset from InstructPix2Pix [11], which provides a modification text and a target caption for 700 input captions. We augment this dataset with 15 annotations that cover additional cases. More details about the additional examples can be found in Section C.4 .